[figure: 9.7 cm]

CORRELATED ROUGHNESS EFFECTS IN THE GIANT MAGNETORESISTANCE IN MAGNETIC MULTILAYERS

G. Palasantzas^{a}, J. Barnaś^{b}, and J.Th.M. De Hosson^{a} a Department of Applied Physics and Materials Science Group, University of Groningen, Nijenborgh 4, 9747 AG Groningen, The Netherlands b Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61–614 Poznań, Poland

We study interface roughness effects on the giant magnetoresistance in magnetic multilayers with correlated interfaces having non-zero cross correlation function. Initially we present the formalism that accounts for interface electron scattering due to cross correlation. The roughness in each interface is described in terms of the K-correlation model and is characterized by the roughness exponent H (0¡H¡1), the correlation length ξ , and the roughness amplitude δ. The magnetoresistance is strongly sensitive to any differences in the parameters H, ξ, and δ of consecutive interfaces. This sensitivity is additionally influenced by the electronic structure of the multilayer (potential steps at the interfaces).

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : Poster

Corresponding author : Dr. G. Palasantzas

Address for correspondence : Department of Applied Physics and Materials Science Group University of Groningen Nijenborgh 4, AG 9747 Groningen, The Netherlands

Email address : G.Palasantzas@Phys.rug.nl